OpenStudy (anonymous):

Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can make corrections as many times as he wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Phillip can receive. Identify this as converging or diverging

OpenStudy (anonymous):

@ganeshie8 @zepdrix

OpenStudy (anonymous):

PLEASE HELP D:

OpenStudy (anonymous):

evidently he has 25 more points to go right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how do i make the formula

OpenStudy (anonymous):

so he can have \(75+25\times \frac{2}{10}\)

OpenStudy (anonymous):

btw does it really say \(\frac{2}{10}\) ?

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

cant you simplify \[\frac{ 2 }{ 10 }\] into \[\frac{ 1 }{ 5 }\]?

OpenStudy (anonymous):

ok on planet earth we usually call that \(\frac{1}{5}\) and you can add up the first time he gets \(25\times \frac{1}{5}=5\) the second time yes of course

OpenStudy (anonymous):

now he has \(80\) points leaving \(20\) hmmm this is not as simple as i thought, but i bet we can still do it

OpenStudy (anonymous):

So what is the formula then? just 75=25 x 1/5?

OpenStudy (anonymous):

75+25*

OpenStudy (anonymous):

the geometric sequence meant to be ar^n-1 u will have 75,80,84 this is the grades the the increasing is 5,4,3.2,2.56 ar^n-1 a= 5 r = 4/5.

OpenStudy (anonymous):

the next time he can get \(20\times \frac{1}{5}=4\) and now he has \(84\) points leaving \(16\) the increase each time is \(25(\frac{1}{5})^n\)

OpenStudy (anonymous):

what @Catch.me said

OpenStudy (anonymous):

and the formula for the sum is \[a(\frac{ 1-r ^{n} }{ 1-r })\] so make that = 25 (the maximum ) \[25= 5(\frac{ 1-(\frac{ 4 }{ 5 })^{n} }{ 1-\frac{ 4 }{ 5 } })\]

OpenStudy (anonymous):

n goes to infinity.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/526ebcb0e4b0c0f84759f7bc @ceol1998 are you with me or i got it wrong??

OpenStudy (anonymous):

i got it right :) thanks

OpenStudy (anonymous):

you are welcome, can you explain to me what converging or diverging mean??